OpenStudy (anonymous):

I need some serious help on this question! SOMEBODY PLEASE HELP!!!!!!!! PLEASE!!!!! A textile manufacturer has daily production costs of C=0.45x^2 - 110x + 10,000 where C is the total cost (in dollars) and x is the number of units produced. How many units should be produced each day to yield a minimum cost? What is the minimum cost?

OpenStudy (anonymous):

Find dC/dx and put dC/dx=0 and find values of x then find \[\frac{ d ^{2}C }{ dx ^{2} } at this point.\]

OpenStudy (anonymous):

Okay, so I'm a bit of an idiot when it comes to math. So where do I start?

OpenStudy (anonymous):

Sorry. Math is just not up my alley.

OpenStudy (anonymous):

\[C=0.45x ^{2}-110x-10,000\] \[\frac{ dC }{dx }=0.92x-110\] \[\frac{ dC }{ dx }=0 gives 0.90x-110=0,x=\frac{ 110 }{0.9 }=\frac{ 1100 }{9 }=122\frac{ 2 }{ 9} =122 \] \[\frac{ d ^{2}C }{dx ^{2} }=0.9\] \[atx=122, \frac{ d ^{2}C }{dx ^{2} }>0 hence C is minimum.\] \[Minimum cost C=0.45\left( 122 \right)^{2}-110*122+10,000\] solve it.

OpenStudy (anonymous):

correction in the beginning write +10,000

OpenStudy (anonymous):

Okay. I think I got the answer now though it took me a while. Thank you Surjithayer. I appreciate the help!!!!!!